OpenStudy (anonymous):

Simplify: -2(7 – a) + 16a

OpenStudy (anonymous):

-14+2a+16a -14+18a

OpenStudy (anonymous):

step one, distributive property \[-2(7-a)+16a=-2\times 7+(-2)\times (-a)+16a=-14+2a+16a\]step two combine like terms \[14+2a+16a=-14+18a\] \[

OpenStudy (anonymous):

satellite73... is step two the simplifyed answer??